Your aunt is planning to invest in a bank CD that will pay 7.5 percent interest semiannually. If she has $12,000 to invest, how much will she have at the end of four years? (Round your answer to the nearest penny.)   
Answer #1

We use the formula:
A=P(1+r/2)^2n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=12000*(1+0.075/2)^(2*4)

=12000*1.34247078

=$16109.65(Approx).